# Locate, Assign, Refine: Taming Customized Image Inpainting with Text-Subject Guidance

### v1

@@ Abstract
Prior studies have made significant progress in image inpainting guided by either text or subject image. However, the research on editing with their combined guidance is still in the early stages. To tackle this challenge, we present LAR-Gen, a novel approach for image inpainting that enables seamless inpainting of masked scene images, incorporating both the textual prompts and specified subjects. Our approach adopts a coarse-to-fine manner to ensure subject identity preservation and local semantic coherence. The process involves (i) Locate: concatenating the noise with masked scene image to achieve precise regional editing, (ii) Assign: employing decoupled cross-attention mechanism to accommodate multi-modal guidance, and (iii) Refine: using a novel RefineNet to supplement subject details. Additionally, to address the issue of scarce training data, we introduce a novel data construction pipeline. This pipeline extracts substantial pairs of data consisting of local text prompts and corresponding visual instances from a vast image dataset, leveraging publicly available large models. Extensive experiments and varied application scenarios demonstrate the superiority of LAR-Gen in terms of both identity preservation and text semantic consistency. Project page can be found at https://ali-vilab.github.io/largen-page/.
Keywords:
Image inpainting Diffusion model Text-subject-guided
Figure 1: Image inpainting results of our method. Given a quadruplet comprising a scene image, a scene mask which marked in gray, a subject image, and a text prompt, our proposed method can accurately inpaint the masked area within the scene image as specified by the scene mask. It does so according to guidance that may be (a) text-only, (b) subject-only, or (c) a combination of text and subject, all within the same model. Best viewed when zoomed in.
1 Introduction
@@
Thanks to the exhilarating advancements achieved by diffusion models [15, 32, 36, 16], image generation [31, 33, 34, 10, 7] is witnessing an exuberant proliferation and has found its application in various scenarios [54, 20, 6]. Image inpainting, which aims to fill in missing regions of an image based on various guiding information, has emerged as a significant application and is exhibiting particularly attractive characteristics recently. Commonly, plain image inpainting technologies [43, 57, 58] leverage the inherent semantics of the original image to appropriately fill in masked regions. Additionally, some auxiliary information could be incorporated to facilitate customized edits. For example, text-guided image inpainting methods [1, 46] adjust images based on user-provided text descriptions. However, such methods often depend on global descriptions of the image, potentially limiting the fidelity of local semantic correspondence within the masked areas. Conversely, subject-guided image inpainting approaches [8, 49, 55] reconstruct the masked region using a subject image. Despite this, the prevalent use of a collage strategy [8, 55] can result in issues resembling a simple copy-paste effect. Besides, both text-only and image-only guided methods lack the precision required for fine control over the reconstruction of missing regions.
@@
To address the aforementioned drawbacks, we first present text-subject-guided image inpainting, a novel task that seamlessly integrates an arbitrary customized object into the desired location within a scene image, and allows for auxiliary text prompt to achieve fine-grained control. This task is elaborately designed to process with a maximum input of quadruplet, including a scene image, a scene mask, a subject image, and a local textual description, adhering to the following distinct principles:
@@
  • The background area of the scene image should remain unchanged, focusing solely on the editing of the specific masked area.
@@
  • The subject identity and details should be preserved as much as possible.
@@
  • The inpainted content should semantically correspond to the local text.
@@
  • The reconstructed image should achieve high-quality and seamless integration within the scene.
@@
We propose a tuning-free method, termed as LAR-Gen, which follows a “Locate, Assign, Refine” pipeline to achieve the above objectives and enable creative Generation. Specifically, the masked scene image is first encoded into latent space and concatenated with the noise input, along with the mask image. This stage compels the model to seamlessly inpaint the masked region while keeping the background unaltered. Then, a decoupled cross-attention mechanism is employed to effectively guide the diffusion process under the joint control of the text and the subject, ensuring that the guidance process conforms to the semantics of the local description and the coarse-grained subject reference. Finally, an auxiliary U-Net [37] termed RefineNet is introduced to supplement subject details. It encodes the noisy subject image at each sampling step and injects the detail features into the main branch at multi-scale self-attention layers, facilitating the subject detail preservation. To address the scarcity of public datasets providing subject images paired with localized textual prompts for the text-subject-guided image inpainting task, we introduce an innovative data construction strategy for assembling the necessary quadruple data. The pipeline operates by automatically categorizing [56], detecting [30], instance segmenting [23], and visual captioning [29], thereby producing region-level quadruples comprising a scene image, a scene mask, a subject image and a text prompt. With the help of the collected quadruple data and the tailored pipeline, LAR-Gen not only excels in text-subject-guided image inpainting task, but also supports text-only and image-only guided inpainting within the same framework, as illustrated in Fig. 1. This indicates that our LAR-Gen is a unified framework for image inpainting and can be employed in various scenarios. Extensive experiments demonstrate the superiority of LAR-Gen in terms of both subject identity consistency and text semantic consistency.
2 Related Work
@@
Text-to-image Generation. In the past few years, diffusion-based image generation [15, 40, 41, 32, 25, 5] has emerged as a burgeoning research trend. This is due to the diffusion models’ superior ability to generate high-quality images compared with GANs [14, 21, 22] and auto-regressive models [11, 12]. Large-scale diffusion models have been proven to be excellent starting points for various downstream tasks, such as image inpainting [1, 48, 46], image super-resolution [44, 28, 50], and video generation [19, 17]. Our LAR-Gen builds upon Stable Diffusion [36, 2, 3] to fully leverage its power on generating a high-fidelity image.
@@
Subject-driven Image Generation. Subject-driven image generation [53, 18, 26, 45] aims to generate images conditioned on a customized subject and a text prompt that describes the context. Existing works can be categorized in terms of the necessity of test-time tuning. One line of them requires one more image to conduct training optimization for specific subjects. Textual Inversion [13] is the first work to inverse a subject into textual representation space. It fixes the U-Net [37] backbone and only tunes the newly added embedding. DreamBooth [38] tunes the entire U-Net together with the registered embedding and introduces an auxiliary loss to prevent generative ability degradation. Custom Diffusion [24] performs parameter updating only at cross-attention layers and is able to combine multiple subjects into one image. While they excel at capturing subject detail, the time-consuming nature of the tuning process renders these methods less efficient than zero-shot approaches, which in turn limits their practical application. Another line of them encodes the subject with a visual encoder, avoiding test-time tuning. InstantBooth [39] and ELITE [47] combine both global and local visual features and inject the fine-grained local features with an extra cross-attention layer to preserve the high-level details of the subject. CustomNet [52] proposes a novel view generation method for the subject, which introduces the camera extrinsic as an extra condition to control the generated view of the subject. IP-Adapter [51] introduces a decoupled cross-attention mechanism, which achieves multi-modal prompts control by weighted mixing of two attention layer outputs. Despite the significant advancements, these methods are still at the early stage of maintaining the subject’s identity. Moreover, their inability to perform localized editing constrains their applicability.
@@
Image Inpainting. Given a masked scene image, image inpainting aims to inpaint the masked region guided by visual or textual prompts. While text-guided image inpainting technology [1, 46] has reached a stage of maturity, subject-driven image inpainting remains in its nascent phase. Recently many works have been proposed to address subject-driven image inpainting. Paint-by-Example [49] replaces the text embedding with global image embedding in the cross-attention layer to inject the subject information into the diffusion process. ObjectStitch [42] introduces a content adapter to mapping the image embedding to textual embedding space and applies the mask to constrain the denoising area, thereby achieving image conditional editing. AnyDoor [8] and PhD [55] apply a ControlNet [54] to control the diffusion process with a subject-scene collage. Although these approaches have achieved significant advancements in subject- or text-conditioned image inpainting, none of them supports multi-modal prompt control. This paper addresses the text-subject-guided image inpainting task, which enables the joint control of visual and textual prompts.
@@
Figure 2: Overall pipeline of LAR-Gen. The whole framework consists of three key mechanisms: the Locate mechanism concatenates the noise with masked scene image and mask to compel the model to seamlessly inpaint the masked region while keeping the background unaltered; the Assign mechanism, indeed a decoupled cross-attention mechansim, achieves text compatible subject guidance on denoising process; the Refine mechanism utilizes a RefineNet to gradually supplement the subject details, facilitating subject identity preservation. The fire icon indicates that the module parameter requires tuning, while the snow icon signifies that no tuning is needed.
3 Method
Overview. Given a scene image $x_{s}\in\mathbb{R}^{H\times W\times 3}$, a binarized mask $m\in\mathbb{R}^{H\times W}$, a subject image $x_{\text{obj}}$ and a text prompt $s$, LAR-Gen aims to inpaint a local region of $x_{s}$ specified by $m$, under the joint control of $x_{\text{obj}}$ and $s$. The inpainted content inside region should preserve the subject identity referring to $x_{\text{obj}}$, and also align with the semantic of text prompt. The overall pipeline of LAR-Gen is depicted in Fig. 2. LAR-Gen reconciles three mechanisms to achieve high-fidelity and seamless composition: Locate, Assign and Refine. We will describe each mechanism in detail next.
3.1 Locate Mechanism
@@
The text-to-image diffusion U-Net [37] receives the latent code of a noisy image $z\in\mathbb{R}^{h\times w\times 4}$ and predicts the noise added to the image. To support local image editing, we concatenate the masked image with noise at channel dimension, following SD-Inpainting [1]. It first encodes the masked scene image, formulated as:
$z_{s}=\text{Enc}(x_{s}\odot(1-m)),$ (1)
@@
$\hat{z}^{t}=z_{s}^{t}\odot(1-m^{*})+{z}^{t}\odot m^{*}$ (3)
where ${z}^{t}$ and $z_{s}^{t}$ denote the model output and the noisy latent code of scene image at sampling step $t$, respectively.
Training Data. The standard inpainting method trains the model with (image, prompt) pair data, where the prompt is a global description of image. However, we find that global caption may ignore the target subject or fail to describe it in detail. This leads to semantic misalignment on our training stage, and sometimes fail to generate subject. To address this, we propose a novel data process method to collect the needed quadruplets. Fig. 3 presents several examples to demonstrate the distinctions between global captions and region-specific captions. We will introduce the data collection strategy in Sec. 3.4.
Figure 3: Comparison between global caption and regional caption. Global captions can overlook specific subjects in complex images, while regional captions offer more precise semantic alignment with local visual content.
3.2 Assign Mechanism
While the pursuit of image inpainting conditioned independently on text descriptions or reference images has garnered considerable attention in recent years, mixing the two type of conditions presents a non-trivial challenge. Inspired by IP-adapter[51], we implement a decoupled cross-attention mechanism to enable compatible multi-modal prompt control. Specifically, given the subject image $x_{\text{obj}}$, we adopt a projection module $f$ attached to CLIP image encoder to encode the subject image to a sequence of features $c_{i}$ with length $N$. It is formulated as:
$c_{i}=f(\text{CLIP}(x_{\text{obj}})),$ (4)
In our experiments, we adopt IP-Adapter-Plus[51] as the projection module $f$ and freeze both CLIP encoder and the projection module during training. The feature sequence length $N$ is 16. Subsequently, two extra parameters $\mathbf{W}_{k}^{{}^{\prime}}$, $\mathbf{W}_{v}^{{}^{\prime}}$ are introduced to insert image feature to cross-attention layer:
$\mathbf{K}^{{}^{\prime}}=c_{i}\mathbf{W}_{k}^{{}^{\prime}},\mathbf{V}^{{}^{\prime}}=c_{i}\mathbf{W}_{v}^{{}^{\prime}},$ (5)
@@ 3.2 Assign Mechanism
$\mathbf{Z}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}})\mathbf{V},$ (6)
$\mathbf{Z}^{{}^{\prime}}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{(K^{{}^{\prime}})}^{\top}}{\sqrt{d}})\mathbf{V^{{}^{\prime}}},$ (7)
@@
$\mathbf{Z}^{\text{out}}=\mathbf{Z}+\beta\mathbf{Z}^{{}^{\prime}},$ (8)
where $\beta$ is a hyper-parameter used to modulate the image control strength.
3.3 Refine Mechanism
To achieve text-subject-compatible guidance, the hyper-parameter $\beta$ in the decoupled cross-attention layer is generally set to a small value to ensure text consistency. However, this compromises image control and may result in the loss of subject details. To counteract this, we introduce an auxiliary U-Net, termed RefineNet, which enhances the object’s details even when $\beta$ is small. The RefineNet performs denoising process on the noisy subject image, which is modulated with the same level of Gaussian noise. It utilizes a detail encoder, comprised of a CLIP image encoder and an MLP, to encode the subject’s patch features:
$c_{i}^{{}^{\prime}}=\text{MLP}(\text{CLIP}(x_{\text{obj}}))$ (9)

### v2

# Locate, Assign, Refine: Taming Customized Promptable Image Inpainting

@@ Abstract.
Prior studies have made significant progress in image inpainting guided by either text description or subject image. However, the research on inpainting with flexible guidance or control, i.e., text-only, image-only, and their combination, is still in the early stage. Therefore, in this paper, we introduce the multimodal promptable image inpainting project: a new task model, and data for taming customized image inpainting. We propose LAR-Gen, a novel approach for image inpainting that enables seamless inpainting of specific region in images corresponding to the mask prompt, incorporating both the text prompt and image prompt. Our LAR-Gen adopts a coarse-to-fine manner to ensure the context consistency of source image, subject identity consistency, local semantic consistency to the text description, and smoothness consistency. It consists of three mechanisms: (i) Locate mechanism: concatenating the noise with masked scene image to achieve precise regional editing, (ii) Assign mechanism: employing decoupled cross-attention mechanism to accommodate multi-modal guidance, and (iii) Refine mechanism: using a novel RefineNet to supplement subject details. Additionally, to address the issue of scarce training data, we introduce a novel data engine to automatically extract substantial pairs of data consisting of local text prompts and corresponding visual instances from a vast image data, leveraging publicly available pre-trained large models. Extensive experiments and various application scenarios demonstrate the superiority of LAR-Gen in terms of both identity preservation and text semantic consistency. Refer to caption Figure 1. We build a unified framework for taming customized promptable image inpainting, consisting three interconnected components: a promptable image inpainting task, a Locate, Assign, Refine for Generation model (LAR-Gen), and a data engine for constructing the data for training LAR-Gen.
1. Introduction
@@
Thanks to the exhilarating advancements achieved by diffusion models (Ho et al., 2020; Nichol et al., 2022; Rombach et al., 2022; Ho and Salimans, 2021), image generation (Midjourney, 2023; OpenAI, 2023; Cloud, 2023; Chen et al., 2024b) is witnessing an exuberant proliferation and has found its application in various scenarios (Zhang et al., 2023c; Jiang et al., 2024; Brooks et al., 2023). Image inpainting, which aims to fill in missing regions of an image using various guiding inputs, has become as a significant application and recently demonstrated particularly attractive characteristics. Commonly, plain image inpainting technologies (Suvorov et al., 2022; Zhao et al., 2021; Zhu et al., 2023) leverage the inherent semantics of the original image to appropriately fill in masked regions. Additionally, some auxiliary information could be incorporated to facilitate customized edits. For example, text-guided image inpainting methods (AI, 2022a; Wang et al., 2023) adjust images based on user-provided text prompt, as shown in the top of LABEL:fig:teaser. However, such methods often depend on global descriptions of the image, potentially limiting the fidelity of local semantic correspondence within the masked areas. Conversely, subject-guided image inpainting methods (Chen et al., 2024a; Yang et al., 2023; Zhang et al., 2023b) reconstruct the masked region based on the referenced image prompt, as shown in the middle of LABEL:fig:teaser. Despite this, the prevalent use of a collage strategy (Chen et al., 2024a; Zhang et al., 2023b) can result in issues resembling a simple copy-paste effect. Besides, both text-only and image-only guided methods lack the precision required for fine control over the reconstruction of missing regions.
Therefore, in this paper, our goal is to build a unified framework for taming customized promptable image inpainting, which means that we seek to develop a multimodal promptable model and train it on automatically constructed data for text and image flexible guided image inpainting, as shown in LABEL:fig:teaser. It hingles on three components: task, model, and data, as shown in Fig. 1.
@@
TASK (Section 3). As shown in Fig. 1 (a), we first introduce a new task that seamlessly integrates an arbitrary customized object (i.e., image prompt) into the desired location (i.e., mask prompt) within a source image, and allows for auxiliary text prompt to achieve fine-grained control. This task is elaborately designed to accept a maximum input of quadruplet, including a source image, a mask prompt, an image prompt containing a subject, and a text prompt, adhering to the following distinct principles:
@@
  • Context consistency. The background area of the source image should remain unchanged, focusing solely on the editing in the specific area of mask prompt.
@@
  • ID consistency. The subject identity and details from image prompt should be preserved as much as possible.
@@
  • Semantic consistency. The inpainted content should semantically correspond to the text prompt.
@@
  • Smoothness consistency. The reconstructed image should achieve seamless integration with the context of the source image, and maintain high quality.
@@
MODEL (Section 4). As shown in Fig. 1 (b), we propose a unified promptable image inpainting model, termed as LAR-Gen, which follows a “Locate, Assign, Refine” pipeline to achieve the above objectives and enable creative Generation. Specifically, the source image is first encoded into latent space and concatenated with the noise input, along with the mask prompt. This stage compels the model to seamlessly inpaint the masked region while keeping the background unaltered. Then, a decoupled cross-attention mechanism is designed to effectively guide the diffusion process under the joint control of the text prompt and the image prompt, ensuring that the guidance process conforms to the semantics of the local textual description and the coarse-grained subject reference. Finally, an auxiliary U-Net (Ronneberger et al., 2015) termed RefineNet is introduced to supplement subject details. It encodes the noisy subject image at each sampling step and injects the detail features into the main branch at multi-scale self-attention layers, facilitating the subject detail preservation. With the help of the collected quadruple data (details in Section 5) and the tailored pipeline, LAR-Gen not only excels in text-subject-guided image inpainting task, but also supports text-only and image-only guided inpainting within the same framework, as illustrated in LABEL:fig:teaser. This indicates that our LAR-Gen is a unified framework for image inpainting and can be employed in various scenarios. Besides, extensive experiments demonstrate the superiority of LAR-Gen in terms of both subject identity consistency and text semantic consistency.
DATA (Section 5). Fig. 1 (c), to address the scarcity of public datasets that provide subject images (i.e., image prompt) paired with localized textual descriptions (i.e., text prompt) for multimodal promptable image inpainting, we introduce an innovative data engine to automatically assemble the necessary quadruple data. The pipeline operates by automatically tagging (Zhang et al., 2024), localizing (Liu et al., 2025), instance segmenting (Kirillov et al., 2023), and visual captioning (Liu et al., 2023), thereby producing region-level quadruples comprising a source image, a mask prompt, an image prompt and a text prompt.
2. Related Work
2.1. Text-to-image Generation
@@
In the past few years, diffusion-based image generation (Ho et al., 2020; Song et al., 2021a, b; Nichol et al., 2022; Li et al., 2023; Avrahami et al., 2022) has emerged as a burgeoning research trend. This is due to the diffusion models’ superior ability to generate high-quality images compared with GANs (Goodfellow et al., 2020; Karras et al., 2019) and auto-regressive models (Ding et al., 2021; Esser et al., 2020). Large-scale diffusion models have been proven to be excellent starting points for various downstream tasks, such as image inpainting (AI, 2022a; Xie et al., 2023; Wang et al., 2023), and image super-resolution (Wang et al., 2024b; Lin et al., 2025; Yang et al., 2025). Our LAR-Gen builds upon Stable Diffusion (Rombach et al., 2022; AI, 2022b) to fully leverage its power on generating a high-fidelity image.
2.2. Subject-driven Image Generation
@@
Subject-driven image generation (Zhang et al., 2023a; Wang et al., 2024a) aims to generate images conditioned on a customized subject and a text prompt that describes the context. Existing works can be categorized in terms of the necessity of test-time tuning. One line of them requires one more image to conduct training optimization for specific subjects. Textual Inversion (Gal et al., 2022) is the first work to inverse a subject into textual representation space. It fixes the U-Net (Ronneberger et al., 2015) backbone and only tunes the newly added embedding. DreamBooth (Ruiz et al., 2023) tunes the entire U-Net together with the registered embedding and introduces an auxiliary loss to prevent generative ability degradation. Custom Diffusion (Kumari et al., 2022) performs parameter updating only at cross-attention layers and is able to combine multiple subjects into one image. While they excel at capturing subject detail, the time-consuming nature of the tuning process renders these methods less efficient than zero-shot approaches, which in turn limits their practical application. Another line of them encodes the subject with a visual encoder, avoiding test-time tuning. InstantBooth (Shi et al., 2024) and ELITE (Wei et al., 2023) combine both global and local visual features and inject the fine-grained local features with an extra cross-attention layer to preserve the high-level details of the subject. CustomNet (Yuan et al., 2023) proposes a novel view generation method for the subject, which introduces the camera extrinsic as an extra condition to control the generated view of the subject. IP-Adapter (Ye et al., 2023) introduces a decoupled cross-attention mechanism, which achieves multi-modal prompts control by weighted mixing of two attention layer outputs. Despite the significant advancements, these methods are still at the early stage of maintaining the subject’s identity. Moreover, their inability to perform localized editing constrains their applicability.
2.3. Image Inpainting
@@
Given a masked scene image, image inpainting aims to inpaint the masked region guided by visual or textual prompts. While text-guided image inpainting technology (AI, 2022a; Wang et al., 2023) has reached a stage of maturity, subject-driven image inpainting remains in its nascent phase. Recently many works have been proposed to address subject-driven image inpainting. Paint-by-Exa (Yang et al., 2023) replaces the text embedding with global image embedding in the cross-attention layer to inject the subject information into the diffusion process. ObjectStitch (Song et al., 2023) introduces a content adapter to mapping the image embedding to textual embedding space and applies the mask to constrain the denoising area, thereby achieving image conditional editing. AnyDoor (Chen et al., 2024a) and PhD (Zhang et al., 2023b) apply a ControlNet (Zhang et al., 2023c) to control the diffusion process with a subject-scene collage. Although these approaches have achieved significant advancements in subject- or text-conditioned image inpainting, none of them supports multi-modal prompt control. This paper addresses the text-subject-guided image inpainting task, which enables the joint control of visual and textual prompts.
3. Promptable Image Inpainting Task
The new promptable image inpainting task is defined as the source image is edited based on the multimodal prompts, i.e., mask prompt, image prompt, and text prompt, as shown in Fig. 1 (a). Specifically, given a source image $x_{s}\in\mathbb{R}^{H\times W\times 3}$, a mask prompt $m\in\mathbb{R}^{H\times W}$, an image prompt $x_{\text{obj}}$ and a text prompt $s$, LAR-Gen aims to inpaint a local region of $x_{s}$ specified by $m$, under the control of $x_{\text{obj}}$ and $s$ separately or jointly. The inpainted content inside region should remain contextual area of the source iamge unchanged (i.e., context consistency), preserve the subject identity referring to image prompt (i.e., ID consistency), align with the semantic of text prompt (i.e., semantic consistency), and achieve high-quality and seamless integration with the source image (i.e., smoothness consistency).
4. LAR-Gen: Promptable Image Inpainting Model
@@
The overall pipeline of LAR-Gen is depicted in Fig. 2, reconciling three mechanisms to achieve high-fidelity and seamless composition: Locate, Assign and Refine. a) Locate mechanism concatenates the noise with source image and mask prompt to compel the model to seamlessly inpaint the masked region while keeping the background unaltered. b) Assign mechanism, indeed a decoupled cross-attention mechanism, achieves text compatible subject guidance on denoising process. c) Refine mechanism gradually supplements the subject details, facilitating subject identity preservation.
4.1. Locate Mechanism
@@
The text-to-image diffusion U-Net (Ronneberger et al., 2015) receives the latent code of a noisy image $z\in\mathbb{R}^{h\times w\times 4}$ and predicts the noise added to the image. To support local image editing, we concatenate the masked image with noise at channel dimension, following SD-Inpainting (AI, 2022a). It encodes the masked source image as:
(1) $z_{s}=\text{Enc}(x_{s}\odot(1-m)),$
@@
(3) $\hat{z}^{t}=z_{s}^{t}\odot(1-m^{*})+{z}^{t}\odot m^{*}$
where ${z}^{t}$ and $z_{s}^{t}$ denote the model output and the noisy latent code of source image at sampling step $t$, respectively.
Figure 2. Overall pipeline of LAR-Gen.
4.2. Assign Mechanism
While the pursuit of image inpainting conditioned independently on text descriptions or reference images has garnered considerable attention in recent years, mixing the two type of conditions presents a non-trivial challenge. Inspired by IP-adapter(Ye et al., 2023), we implement a decoupled cross-attention mechanism to enable compatible multi-modal prompt control. Specifically, given the subject image prompt $x_{\text{obj}}$, we adopt a projection module $f$ attached to CLIP image encoder to encode the subject image prompt to a sequence of features $c_{i}$ with length $N$. It is formulated as:
(4) $c_{i}=f(\text{CLIP}(x_{\text{obj}})),$
In our experiments, we adopt IP-Adapter-Plus(Ye et al., 2023) as the projection module $f$ and freeze both CLIP encoder and the projection module during training. The feature sequence length $N$ is 16. Subsequently, two extra parameters $\mathbf{W}_{k}^{{}^{\prime}}$, $\mathbf{W}_{v}^{{}^{\prime}}$ are introduced to insert image feature to cross-attention layer:
(5) $\mathbf{K}^{{}^{\prime}}=c_{i}\mathbf{W}_{k}^{{}^{\prime}},\mathbf{V}^{{}^{\prime}}=c_{i}\mathbf{W}_{v}^{{}^{\prime}},$
@@ 4.2. Assign Mechanism
(6) $\mathbf{Z}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}})\mathbf{V},$ $\mathbf{Z}^{{}^{\prime}}=\text{Softmax}(\frac{\mathbf{Q}\mathbf{(K^{{}^{\prime}})}^{\top}}{\sqrt{d}})\mathbf{V^{{}^{\prime}}},$
@@
(7) $\mathbf{Z}^{\text{out}}=\mathbf{Z}+\beta\mathbf{Z}^{{}^{\prime}},$
where $\beta$ is a hyper-parameter to modulate image control strength.
4.3. Refine Mechanism
To achieve text-subject-compatible guidance, the hyper-parameter $\beta$ in the decoupled cross-attention layer is generally set to a small value to ensure text consistency. However, this compromises image control and may result in the loss of subject details. To counteract this, we introduce an auxiliary U-Net, termed RefineNet, which enhances the object’s details even when $\beta$ is small. The RefineNet performs denoising process on the noisy subject image prompt, which is modulated with the same level of Gaussian noise. It utilizes a detail encoder, comprised of a CLIP image encoder and an MLP, to encode the subject’s patch features:
(8) $c_{i}^{{}^{\prime}}=\text{MLP}(\text{CLIP}(x_{\text{obj}}))$